# On the adjoint Selmer groups of semi-stable elliptic curves and Flach’s zeta elements

### v4

@@ 3. Verifying the hypotheses for Stark systems
We review the running hypotheses of Stark systems (of core rank one) in [MR16, §$\S$4] and verify them for our setting.
@@ 3.1. Running hypotheses
We list the assumptions to proceed the Stark system argument [MR16, §$\S$4]:
@@
    $I_{q}=0$ for $q\in\mathcal{P}^{\mathrm{Flach},(\ell)}_{k}$ when the coefficient ring is $\mathbb{Z}/p^{k}\mathbb{Z}$.
@@
(H.1) follows from Lemma 3.3 and [Rub00, §$\S$3.6].
@@
This choice of $\tau$ is possible and satisfies (H.2) for $\mathrm{Sym}^{2}(T_{p}E)$ as explained in [Rub00, §$\S$3.6]. The choice of $\tau$ determines the type of auxiliary primes when we apply the Chebotarev density theorem as in Definition 2.3. In the case of Heegner points, $\tau$ is chosen to be the complex conjugation. In [Fla92], $\tau$ is chosen to be the complex conjugation again. For Kato’s Euler systems, $\tau$ is chosen to be a unipotent element as explained in [Rub00, Prop. 3.5.8]. We also recommend the reader to read [Wil95, Intro.] for his change of auxiliary primes.
@@
(H.3) follows from [Rub00, §$\S$3.6]. See also [Fla92, Lem. 1.2] with [Fla95, Rem. 1 in §$\S$4], and [DDT97, Lem. 2.48] when $p>5$.
@@
Suppose that we are working over the $\mathbb{Z}/p^{k}\mathbb{Z}$-coefficients. For $q\in\mathcal{P}^{\mathrm{Flach},(\ell)}_{k}$, $I_{q}$ is defined by the ideal of $\mathbb{Z}/p^{k}\mathbb{Z}$ generated by $(1-\alpha^{2}_{q})\cdot(1-q)\cdot(1-\beta^{2}_{q})$ and $(1-q)$. This means that $I_{q}=(1-q)=(0)$, so we are done.
@@ 5.1. Basic setup for Stark systems
throughout this section.
@@
Following [MR16, §$\S$6], recall that
@@
[…]
@@ 5.2. Definition and properties of Stark systems
The $\mathbb{Z}/p^{k}\mathbb{Z}$-module of the (rank one) Stark system $\mathbf{SS}_{1}(\mathrm{Sym}^{2}(E[p^{k}]))=\mathbf{SS}_{1}(\mathrm{Sym}^{2}(E[p^{k}]),\mathcal{F}^{\ell}_{\mathrm{BK}},\mathcal{P}^{\mathrm{Flach},(\ell)}_{k})$ for Selmer data $(\mathrm{Sym}^{2}(E[p^{k}]),\mathcal{F}^{\ell}_{\mathrm{BK}},\mathcal{P}^{\mathrm{Flach},(\ell)}_{k})$ is defined to be the inverse limit
@@ 5.3. The key computation
We keep the choices of $\ell$ and $n$ in (5.1) from now on. We explicitly compute
@@
where $c(\ell)$ and $c(q_{i})$’s are Flach’s zeta elements in Theorem 2.2 and $\psi_{q_{i}}:\mathrm{H}^{1}_{/f}(\mathbb{Q}_{q_{i}},\mathrm{Sym}^{2}(E[p^{k}]))\simeq\mathbb{Z}/p^{k}\mathbb{Z}$ is the map in Lemma 4.1. Here we use the same notation $c(\ell)$ and $c(q_{i})$ for their image in $\mathrm{H}^{1}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))$. Note that $c(\ell)$ and all $c(q_{i})$’s are linearly independent due to their local conditions in Theorem 2.2. We first apply [MR16, Prop. A.1] to our setting. Consider exact sequence
@@
Then there exists a unique map
@@ 5.4. A Stark system of Gauss sum type and its extension
We keep the choices of $\ell$ and $n$ in (5.1) as before. Define
$\epsilon^{\mathrm{Flach}?,\ell}_{n}:=(-1)^{\frac{s(s+1)}{2}}\cdot c(\ell)\wedge\bigwedge^{s}_{i=1}c(q_{i})\otimes\bigwedge^{s}_{i=1}\psi_{q_{i}}\in Y_{n}$
@@
$\Psi_{n,1}(\epsilon^{\mathrm{Flach}?,\ell}_{n})=\left(\prod_{i=1}^{s}\psi_{q_{i}}\circ\mathrm{loc}^{s}_{q_{i}}(c(q_{i}))\right)\cdot c(\ell).$
This formula follows from the iteration of (5.2) via (5.3).
@@
$\epsilon^{\mathrm{Flach}?,\ell}_{m}:=\Psi_{n,m}(\epsilon^{\mathrm{Flach}?,\ell}_{n}).$
By Definition 5.3, $\left\{\epsilon^{\mathrm{Flach}?,\ell}_{m}\in Y_{m}:m|n\right\}$ forms a finite Stark system of core rank one, i.e. a Stark system for Selmer data $(\mathrm{Sym}^{2}(E[p^{k}]),\mathcal{F}^{\ell}_{\mathrm{BK}},\mathcal{P}^{n})$ where $\mathcal{P}^{n}$ is the (finite!) set of the primes dividing $n$.
Let $m$ is a divisor of $n$ and $G_{m}=\otimes_{q|m}\mathrm{Gal}(\mathbb{Q}(\zeta^{(p)}_{q})/\mathbb{Q})$ where $\mathbb{Q}(\zeta^{(p)}_{q})$ is the maximal $p$-subextension of $\mathbb{Q}$ in $\mathbb{Q}(\zeta_{q})$. Recall the map in [MR16, §$\S$12]
$\Pi_{m}:Y_{m}\to\mathrm{Sel}_{\mathcal{F}^{\ell}_{\mathrm{BK}}(m)}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))\otimes G_{m}$
@@
$\left\{(-1)^{\nu(m)}\cdot\Pi_{m}(\epsilon^{\mathrm{Flach}?,\ell}_{m}):m|n\right\}$
forms a finite Kolyvagin system for $\left(\mathrm{Sym}^{2}(E[p^{k}]),\mathcal{F}^{\ell}_{\mathrm{BK}},\mathcal{P}^{n}\right)$. Thanks to the core rank one property recalled in §$\S$3.7, the finite Kolyvagin system extends to the rank one Kolyvagin system $\boldsymbol{\kappa}^{\mathrm{Flach}?,\ell}$ for $\left(\mathrm{Sym}^{2}(E[p^{k}]),\mathcal{F}^{\ell}_{\mathrm{BK}},\mathcal{P}^{\mathrm{Flach},(\ell)}_{k}\right)$ as in the case of Gauss sum Kolyvagin systems [MR04, Rem. 6.3.3].
@@
If $n\in\mathcal{P}^{\mathrm{Flach},(\ell)}_{k}$ with $\mathrm{Sel}_{\mathcal{F}_{\mathrm{BK},\ell n}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{k}]))=0$, then $\kappa^{\mathrm{Flach}?,\ell}_{n}$ is uniquely determined by $\kappa^{\mathrm{Flach}?,\ell}_{m}$ for $m$ properly dividing $n$ [MR04, Rem. 3.1.9] where $\mathcal{F}_{\mathrm{BK},\ell n}$ is the $\ell n$-strict Bloch–Kato Selmer structure. From this point of view, the extension of a finite Kolyvagin system is not a miracle and it shows the strength of the rigidity of Kolyvagin systems.
The modules of Stark and Kolyvagin systems over $\mathbb{Z}/p^{k}\mathbb{Z}$ are isomorphic as free $\mathbb{Z}/p^{k}\mathbb{Z}$-modules of rank one (Theorem 5.4 and [MR16, Rem. 11.5 and Thm. 12.4]). By using this isomorphism, the finite Stark system also extends to the rank one Stark system $\boldsymbol{\epsilon}^{\mathrm{Flach}?,\ell}$ for $\left(\mathrm{Sym}^{2}(E[p^{k}]),\mathcal{F}^{\ell}_{\mathrm{BK}},\mathcal{P}^{\mathrm{Flach},(\ell)}_{k}\right)$.
5.5. The primitive normalization
We construct the primitive Stark system $\boldsymbol{\epsilon}^{\mathrm{Flach},\ell}$ from $\boldsymbol{\epsilon}^{\mathrm{Flach}?,\ell}$ and compute the constant multiple in Theorem 1.1.
@@ 5.5. The primitive normalization
$\kappa^{\mathrm{Flach}?,\ell}_{1}$ $=\epsilon^{\mathrm{Flach}?,\ell}_{1}$
@@
where the last isomorphism is non-canonical and follows from [MR16, Cor. 3.5.(i)]. By [MR04, Thm. 4.4.1], we know $\langle\kappa^{\mathrm{Flach}?,\ell}_{1}\rangle\subseteq\mathbb{Z}/p^{k}\mathbb{Z}$ in the above decomposition. Thus, we have equality
(5.5) $\mathrm{ord}_{p}(c(\ell))+\mathrm{length}_{\mathbb{Z}_{p}}\left(\dfrac{\mathrm{H}^{1}_{/f}(\mathbb{Q}_{\ell},\mathrm{Sym}^{2}(E[p^{k}]))}{\mathrm{loc}^{s}_{p}\left(\mathrm{Sel}_{\mathcal{F}^{\ell}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))\right)}\right)^{\vee}=\mathrm{ord}_{p}(\mathrm{loc}^{s}_{p}(c(\ell)))$
@@
Write $\lambda=\mathrm{length}_{\mathbb{Z}_{p}}\mathrm{Sel}_{\mathcal{F}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{k}]))$ for convenience. Consider the tautological bound
$\mathrm{length}_{\mathbb{Z}_{p}}\mathrm{Sel}_{\mathcal{F}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{k}]))=\lambda=\mathrm{ord}_{p}\left(\dfrac{p^{\lambda}}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}\cdot\mathrm{loc}^{s}_{\ell}(c(\ell))\right)$
where $\psi_{\ell}:\mathrm{H}^{1}_{/f}(\mathbb{Q}_{\ell},\mathrm{Sym}^{2}(E[p^{k}]))\simeq\mathbb{Z}/p^{k}\mathbb{Z}$ is the map in Lemma 4.1. By (5.4) and (5.5), we also obtain the exact bound the $\ell$-strict one
(5.6) $\mathrm{length}_{\mathbb{Z}_{p}}\mathrm{Sel}_{\mathcal{F}_{\mathrm{BK},\ell}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{k}]))=\mathrm{ord}_{p}\left(\dfrac{p^{\lambda}}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}\cdot c(\ell)\right).$
We apply Theorem 5.5 to $\boldsymbol{\epsilon}^{\mathrm{Flach}?,\ell}$ and compare it with (5.6). Then we obtain
(5.7) $\begin{split}&\mathrm{length}_{\mathbb{Z}_{p}}\mathrm{Sel}_{\mathcal{F}_{\mathrm{BK},\ell}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{k}]))\\ &=\mathrm{ord}_{p}\left(\left(\prod_{i=1}^{s}\psi_{q_{i}}\circ\mathrm{loc}^{s}_{q_{i}}(c(q_{i}))\right)\cdot c(\ell)\right)+\mathrm{ord}_{p}\left(\dfrac{p^{\lambda}}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}\right)-\mathrm{ord}_{p}\left(\prod_{i=1}^{s}\psi_{q_{i}}\circ\mathrm{loc}^{s}_{q_{i}}(c(q_{i}))\right),\end{split}$
so the minimal valuation of the whole Stark system $\boldsymbol{\epsilon}^{\mathrm{Flach}?,\ell}$ is
$\mathrm{ord}_{p}\left(\dfrac{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}{p^{\lambda}}\cdot\prod_{i=1}^{s}\psi_{q_{i}}\circ\mathrm{loc}^{s}_{q_{i}}(c(q_{i}))\right).$
Thus, we define the primitive Flach–Stark system by
$\boldsymbol{\epsilon}^{\mathrm{Flach},\ell}=\dfrac{p^{\lambda}}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}\cdot\left(\prod_{i=1}^{s}\dfrac{1}{\psi_{q_{i}}\circ\mathrm{loc}^{s}_{q_{i}}(c(q_{i}))}\right)\cdot\boldsymbol{\epsilon}^{\mathrm{Flach}?,\ell}$
in the module of Stark systems, and we have explicit formulas
$\epsilon^{\mathrm{Flach},\ell}_{1}$ $=\dfrac{p^{\lambda}}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}\cdot c(\ell),$
$\epsilon^{\mathrm{Flach},\ell}_{n^{\prime}}$ $=\dfrac{(-1)^{\frac{s^{\prime}(s^{\prime}+1)}{2}}\cdot p^{\lambda}}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}\cdot\left(\prod_{i=1}^{s^{\prime}}\dfrac{1}{\psi_{q^{\prime}_{i}}\circ\mathrm{loc}^{s}_{q^{\prime}_{i}}(c(q^{\prime}_{i}))}\right)\cdot c(\ell)\wedge\bigwedge^{s^{\prime}}_{i=1}c(q^{\prime}_{i})\otimes\bigwedge^{s^{\prime}}_{i=1}\psi_{q^{\prime}_{i}}$
for every $n^{\prime}=q^{\prime}_{1}\cdot q^{\prime}_{2}\cdot\cdots\cdot q^{\prime}_{s^{\prime}}\in\mathcal{N}^{\mathrm{Flach},(\ell)}_{k}$ with $\nu(n^{\prime})=s^{\prime}\geq 1$. Note that our choice of $n\in\mathcal{N}^{\mathrm{Flach},(\ell)}_{k}$ in (5.1) was arbitrary (Remark 5.1).
In particular, the constant multiple in Theorem 1.1 is $\dfrac{p^{\lambda}}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}=\dfrac{p^{\lambda}}{\mathrm{deg}(\phi)}$ by construction and by the explicit reciprocity law (2.1). Therefore, the Bloch–Kato conjecture (Corollary 1.4.(2)) is equivalent to showing that this constant multiple is a $p$-adic unit.
@@
Although each $\dfrac{c(q^{\prime}_{i})}{\psi_{q^{\prime}_{i}}\circ\mathrm{loc}^{s}_{q^{\prime}_{i}}(c(q^{\prime}_{i}))}$ makes sense if and only if $\mathrm{loc}^{s}_{q^{\prime}_{i}}$ is surjective, $\epsilon^{\mathrm{Flach},\ell}_{n^{\prime}}$ makes sense without assuming the surjectivity of each $\mathrm{loc}^{s}_{q^{\prime}_{i}}$. In fact, the possible failure of the surjectivity of these localization maps makes the Stark system argument non-vacuous.

### v6

@@ 3. Verifying the hypotheses for Stark systems
We review the running hypotheses of Stark systems (of core rank one) in [MR16, $\S$4] and verify them for our setting.
@@ 3.1. Running hypotheses
We list the assumptions to proceed the Stark system argument [MR16, $\S$4]:
@@
    $I_{q}=0$ for $q\in\mathcal{P}^{\mathrm{Flach},(\ell)}_{k}$ when the coefficient ring is artinian, i.e. $\mathbb{Z}/p^{k}\mathbb{Z}$ in our case.
@@
(H.1) follows from Lemma 3.3 and [Rub00, $\S$3.6].
@@
This choice of $\tau$ is possible and satisfies (H.2) for $\mathrm{Sym}^{2}(T_{p}E)$ as explained in [Rub00, $\S$3.6]. The choice of $\tau$ determines the type of auxiliary primes when we apply the Chebotarev density theorem as in Definition 2.3. In the case of Heegner points, $\tau$ is chosen to be the complex conjugation. In [Fla92], $\tau$ is chosen to be the complex conjugation again. For Kato’s Euler systems, $\tau$ is chosen to be a unipotent element as explained in [Rub00, Prop. 3.5.8]. We also recommend the reader to read [Wil95, Intro.] for his change of auxiliary primes.
@@
(H.3) follows from [Rub00, $\S$3.6]. See also [Fla92, Lem. 1.2] with [Fla95, Rem. 1 in $\S$4], and [DDT97, Lem. 2.48] when $p>5$.
@@
Suppose that we are working over the $\mathbb{Z}/p^{k}\mathbb{Z}$-coefficients. For $q\in\mathcal{P}^{\mathrm{Flach},(\ell)}_{k},$ $I_{q}$ is defined by the ideal of $\mathbb{Z}/p^{k}\mathbb{Z}$ generated by $(1-\alpha^{2}_{q})\cdot(1-q)\cdot(1-\beta^{2}_{q})$ and (1q)(1-q). This means that $I_{q}=(1-q)=(0)$, so we are done.
@@ 5.1. Basic setup for Stark systems
such that $\mathrm{deg}(\phi)^{\nu(n)}=\mathrm{deg}(\phi)^{s}\ll k$ throughout this section.
@@
  1. (1)
    The last condition on $k$ means that we first need to choose large $k$ in order to work with $n\in\mathcal{N}^{\mathrm{Flach},(\ell)}_{k}$ with large $\nu(n)$. This assumption ensures that the result of the key computation in §5.3 is non-vacuous [^6] We deeply thank an anonymous referee for pointing out this condition.. See Remark 5.7.
  2. (2)
@@
Following [MR16, $\S$6], recall that
@@
${\mathrm{Sel}_{\mathcal{F}^{\ell m}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\oplus_{q|m}\mathrm{loc}^{s}_{q}}$${\mathrm{Sel}_{\mathcal{F}^{\ell n}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\oplus_{q|n}\mathrm{loc}^{s}_{q}}$${\bigoplus_{q|m}\mathrm{H}^{1}_{/f}(\mathbb{Q}_{q},\mathrm{Sym}^{2}(E[p^{k}]))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bigoplus_{q|n}\mathrm{H}^{1}_{/f}(\mathbb{Q}_{q},\mathrm{Sym}^{2}(E[p^{k}]))}$
@@ 5.2. Definition and properties of Stark systems
[…]
@@ 5.3. The key computation
We keep the choices of $k$, $\ell$, and $n$ in §5.1 from now on. We explicitly compute
@@
where $c(\ell)$ and $c(q_{i})$’s are Flach’s zeta elements in Theorem 2.2 and $\psi_{q_{i}}:\mathrm{H}^{1}_{/f}(\mathbb{Q}_{q_{i}},\mathrm{Sym}^{2}(E[p^{k}]))\simeq\mathbb{Z}/p^{k}\mathbb{Z}$ is the map in Lemma 4.1. Here we use the same notation $c(\ell)$ and $c(q_{i})$ for their image in $\mathrm{H}^{1}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))$. Note that $c(\ell)$ and all $c(q_{i})$’s are linearly independent due to their local conditions in Theorem 2.2. We first recall a proposition of Mazur–Rubin and apply it to our setting.
Proposition 5.6 (Mazur–Rubin).
Let $R$ be a local principal ideal ring with maximal ideal $\mathfrak{m}$. Suppose that
$0\to N\to M\xrightarrow{\psi}C$
is an exact sequence of finitely generated $R$-modules, with $C$ free of rank one, and $r\geq 1$. Then there exists a unique map
$\widehat{\psi}:\wedge^{r}M\to C\otimes\wedge^{r-1}N$
such that
  1. (1)
    the composition
    $\wedge^{r}M\xrightarrow{\widehat{\psi}}C\otimes\wedge^{r-1}N\to C\otimes\wedge^{r-1}M$
    is given by
    $m_{1}\wedge\cdots\wedge m_{r}\mapsto\sum^{r}_{i=1}(-1)^{i+1}\cdot\psi(m_{i})\otimes(m_{1}\wedge\cdots\wedge m_{i-1}\wedge m_{i+1}\wedge\cdots\wedge m_{r}),$
  2. (2)
    the image of $\widehat{\psi}$ is the image of $\psi(M)\otimes\wedge^{r-1}N\to C\otimes\wedge^{r-1}N$.
If $M$ is free of rank $r$ over $R$, then $\widehat{\psi}$ is an isomorphism if and only if $\psi$ is surjective.
Proof.
See [MR16, Prop. A.1]. ∎
Consider exact sequence
@@
By applying Proposition 5.6 to the above sequence, we obtain a unique map
@@ 5.4. A Stark system of Gauss sum type and its extension
We keep the choices of $k$, $\ell$, and $n$ in §5.1 as before. Define
$\epsilon^{\mathrm{Flach},\ell}_{n}:=(-1)^{\frac{s(s+1)}{2}}\cdot c(\ell)\wedge\bigwedge^{s}_{i=1}c(q_{i})\otimes\bigwedge^{s}_{i=1}\psi_{q_{i}}\in Y_{n}$
@@
[…]
$=\mathrm{deg}(\phi)^{s}\cdot c(\ell)$
This formula follows from the iteration of (5.2) via (5.3) and Proposition 2.4.
Remark 5.7.
Due to our choice of $k$, $\ell$, and $n$ following §5.1, $\mathrm{deg}(\phi)^{s}\cdot c(\ell)$ does not vanish in $\mathrm{Sel}_{\mathcal{F}^{\ell}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))$.
@@
[…]
By Definition 5.3, $\left\{\epsilon^{\mathrm{Flach},\ell}_{m}\in Y_{m}:m|n\right\}$ forms a finite Stark system of core rank one, i.e. a Stark system for Selmer data $(\mathrm{Sym}^{2}(E[p^{k}]),\mathcal{F}^{\ell}_{\mathrm{BK}},\mathcal{P}^{n})$ where $\mathcal{P}^{n}$ is the (finite!) set of the primes dividing $n$.
[…]
(5.4) $\Pi_{m}:Y_{m}\to\mathrm{Sel}_{\mathcal{F}^{\ell}_{\mathrm{BK}}(m)}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))\otimes G_{m}$
@@
$\left\{(-1)^{\nu(m)}\cdot\Pi_{m}(\epsilon^{\mathrm{Flach},\ell}_{m}):m|n\right\}$
[…]
@@
  1. (1)
    The $\mathrm{deg}(\phi)^{\nu(n)}\ll k$ condition in §5.1 is removed in the extended Kolyvagin system $\boldsymbol{\kappa}^{\mathrm{Flach},\ell}$.
  2. (2)
    If $n\in\mathcal{P}^{\mathrm{Flach},(\ell)}_{k}$ with $\mathrm{Sel}_{\mathcal{F}_{\mathrm{BK},\ell n}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{k}]))=0$, then $\kappa^{\mathrm{Flach},\ell}_{n}$ is uniquely determined by $\kappa^{\mathrm{Flach},\ell}_{m}$ for $m$ properly dividing $n$ [MR04, Rem. 3.1.9] where $\mathcal{F}_{\mathrm{BK},\ell n}$ is the $\ell n$-strict Bloch–Kato Selmer structure. From this point of view, the extension of a finite Kolyvagin system is not a miracle and it shows the strength of the rigidity of Kolyvagin systems.
[…]
5.5. The constant multiple
We compare $\boldsymbol{\epsilon}^{\mathrm{Flach},\ell}$ with the primitive one and compute the constant multiple in Theorem 1.1.
@@ 5.5. The constant multiple
$\kappa^{\mathrm{Flach},\ell}_{1}$ $=\epsilon^{\mathrm{Flach},\ell}_{1}$
@@
where the last isomorphism is non-canonical and follows from [MR16, Cor. 3.5.(i)]. By [MR04, Thm. 4.4.1], we know $\langle\kappa^{\mathrm{Flach},\ell}_{1}\rangle\subseteq\mathbb{Z}/p^{k}\mathbb{Z}$ in the above decomposition. Thus, we have equality
(5.6) $\mathrm{ord}_{p}(c(\ell))+\mathrm{length}_{\mathbb{Z}_{p}}\left(\dfrac{\mathrm{H}^{1}_{/f}(\mathbb{Q}_{\ell},\mathrm{Sym}^{2}(E[p^{k}]))}{\mathrm{loc}^{s}_{\ell}\left(\mathrm{Sel}_{\mathcal{F}^{\ell}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{Sym}^{2}(E[p^{k}]))\right)}\right)^{\vee}=\mathrm{ord}_{p}(\mathrm{loc}^{s}_{\ell}(c(\ell)))$
@@
[…]
Now we have
[…]
$=\mathrm{ord}_{p}\left(\dfrac{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}\left(\kappa^{\mathrm{prim},\ell}_{1}\right)}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}\cdot\mathrm{loc}^{s}_{\ell}(c(\ell))\right)$
$=\mathrm{ord}_{p}\left(\dfrac{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}\left(\kappa^{\mathrm{prim},\ell}_{1}\right)}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}\right)\cdot\mathrm{ord}_{p}\left(\mathrm{loc}^{s}_{\ell}(c(\ell))\right)$
$=\mathrm{ord}_{p}\left(\dfrac{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}\left(\kappa^{\mathrm{prim},\ell}_{1}\right)}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}\right)\cdot\mathrm{ord}_{p}\left(\mathrm{deg}(\phi)\right)$
where the first equality follows from the above equality with the global duality argument above, the last equality follows from Proposition 2.4, and $\psi_{\ell}:\mathrm{H}^{1}_{/f}(\mathbb{Q}_{\ell},\mathrm{Sym}^{2}(E[p^{k}]))\simeq\mathbb{Z}/p^{k}\mathbb{Z}$ is the map in Lemma 4.1. In particular, the constant multiple in Theorem 1.1 is
$\dfrac{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}\left(\kappa^{\mathrm{prim},\ell}_{1}\right)}{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}(c(\ell))}=\dfrac{\psi_{\ell}\circ\mathrm{loc}^{s}_{\ell}\left(\kappa^{\mathrm{prim},\ell}_{1}\right)}{\mathrm{deg}(\phi)}=\dfrac{p^{\mathrm{length}_{\mathbb{Z}_{p}}\mathrm{Sel}_{\mathcal{F}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{k}]))}}{\mathrm{deg}(\phi)}$
(up to a $p$-adic unit) by the above computation. Therefore, the Bloch–Kato conjecture (Corollary 1.4.(2)) is equivalent to showing that this constant multiple is a $p$-adic unit.
@@
  1. (1)
    The primitive normalization $\boldsymbol{\epsilon}^{\mathrm{Flach}^{\prime},\ell}$ of $\boldsymbol{\epsilon}^{\mathrm{Flach},\ell}$ is defined by
    $\boldsymbol{\epsilon}^{\mathrm{Flach}^{\prime},\ell}=\dfrac{p^{\mathrm{length}_{\mathbb{Z}_{p}}\mathrm{Sel}_{\mathcal{F}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{k}]))}}{\mathrm{deg}(\phi)}\cdot\boldsymbol{\epsilon}^{\mathrm{Flach},\ell}.$
  2. (2)
    The primitive normalization $\boldsymbol{\kappa}^{\mathrm{Flach}^{\prime},\ell}$ of $\boldsymbol{\kappa}^{\mathrm{Flach},\ell}$ is also defined by
    $\boldsymbol{\kappa}^{\mathrm{Flach}^{\prime},\ell}=\dfrac{p^{\mathrm{length}_{\mathbb{Z}_{p}}\mathrm{Sel}_{\mathcal{F}_{\mathrm{BK}}}(\mathbb{Q},\mathrm{ad}^{0}(E[p^{k}]))}}{\mathrm{deg}(\phi)}\cdot\boldsymbol{\kappa}^{\mathrm{Flach},\ell}$
    and it is equal to $\boldsymbol{\kappa}^{\mathrm{prim},\ell}$ up to a $p$-adic unit.